# This work was supported by project from DOE-CESER

### v1

@@ VI-I Shunt Reactors
Shunt reactors are important reactive management power devices in the power grid. The shunt reactors are crucial in ensuring the voltages in the power network do not rise to dangerously high values causing the insulation in the system to fail. There are: 1) fixed rating shunt reactors either dry type or oil immersed, and; 2) variable shunt reactors mostly oil immersed. The fixed shunt reactors cannot provide regulation while the variable shunt reactors provide regulation capability that is necessary with variable renewable generation like wind and solar [157]. The variable shunt reactors have multiple components including a tap changer and cooling system that are vulnerable to extreme temperatures.
Impact of Climate Change: Like oil cooled transformers, the shunt reactors also have temperature vulnerable components like insulation, windings, tap changer, etc. Reference [158] provides details of the temperature limitations of medium voltage to high voltage oil immersed transformers. Extreme heat (temperatures greater than $40\degree C$) and extreme cold ( temperatures less than $-25\degree C$) ambient conditions are defined as stressed ambient conditions for the oil immersed shunt reactors [158]. There are further details about average temperatures: The monthly average temperature should not exceed $30\degree C$ and annual average should not exceed $20\degree C$ [158]. If the temperatures exceed these ranges, appropriate measures are needed to prevent the impact on reliability and life of the shunt reactor. These temperatures should be maintained during transport and storage [158]. Reference [159] provides details of mechanical failure and microcracks that can develop in dry type of shunt reactors. It shows the mechanism of this failure in extreme cold conditions. This can lead to structural failure and eventual fire hazard. This impact i snot only the failure of the reactor but can easily cascade into a fire risk during polar vortex conditions where the wind speeds are relatively high.
Existing Methods to Measure the Impacts: There are not online measurement methods, however Ref [160] and [161] provide some general guidelines on the performance of the shunt reactors and that could be used to assess the impacts of extreme temperatures on the shunt reactor’s performance or the impact of extreme temperatures.
Existing Methods to Mitigate the Impacts: There are not clear insights on the different methods of mitigating the impacts of extreme temperatures on shunt reactors. Since the size and form factor for some of these high voltage reactors are fairly large it is a challenging problem to shield these from external extreme ambient temperature conditions.
@@ VI-J SVCs and FACTS Devices
SVC (Static Var Compensator) and STATCOM (Static Synchronous Compensator) are shunt devices in the FACTS family. SVCs based on thyristor technology appeared in the field during 1970s. STATCOMs based on GTOs (Gate Turn-off Thyristors) appeared in 1980s but with the help of high power IGBTs (Insulated Gate Bipolar Transistor) in 1990s became commercially available. Both SVC and STATCOM assist in increasing the maximum transfer capability of the power lines. This in turn helps with delaying the the need to building new power lines in face of ever growing power demand on the electric grid [162]. When the space in substations is a constraint. Furthermore, STATCOMs do not have capacitors like SVCs which comes with few advantages in terms of impact due to extreme temperatures.
SVC and STATCOM offer benefits in different way like 1) Voltage control during normal and contingency conditions, 2) Reactive power support during contingencies, 3) Voltage stability, 4) Over voltage reduction during lightly loaded conditions, 5) Damping oscillations, 6) Maximize power transfer capability, and 7) Power quality.
Impact of Climate Change: SVCs face the same point of failures as capacitors and electronics-based devices because of the components within. However, STATCOMs face similar point of failures as electronics-based equipment. Some of the conditions that effect all the outdoor equipment are as follows [163], 1) Ambient temperature, 2) Seismic activity, 3) Wind velocity, 4) Climatic conditions: High humidity, heavy rainfall or snow, sand-storms, ice load, etc., and 5) Pollution: Dust, corrosive atmosphere, etc.
Existing Methods to Measure the Impacts: Existing methods to monitor the impacts is remote monitoring of ambient temperature of the immediate air surrounding the equipment through human machine interface [163].
Existing Methods to Mitigate the Impacts: Similar to frequency changers, these issues can be resolved by placing the equipment in a climate control environment/enclosure. This is especially useful when the climate conditions can change drastically at a specific location.
@@ VI-K Switchgear Panels and Control Panels
Switchgear is a class of equipment that house many different critical components in a substation, switchyard or key locations where multiple paths of the power network interface. There is a class of switchgear that is indoors that is shielded from the environmental conditions to some extent, but there is a class of switchgear that is outdoor and are prone to impacts of environmental conditions. In medium voltage and low voltage systems the outdoor switchgear is often metal enclosed switchgear. These typically provide weather proof housing, however, cannot eliminate the impacts of extreme heat and extreme cold caused by heat domes and polar vortexes or winter storms.
There is also outdoor switchgear is typically located in substations for protecting / managing key substation equipment like transformers, circuit breakers [164]. The radiators and oil pumps are all dedicated low voltage auxiliary components like motors that are managed with their own switchgear these are also called as the control cabinets mounted either on the transformer or near the transformer (21).
@@
Impact of Climate Change: Overheating and insulation failures are the key failure modes for switchgear [165]. The switchgear equipment located in the substation yard is not only subject to weather related ambient conditions but also due to the heat generated by neighboring equipment. Especially during extreme heat conditions, the load/ demand in the system is high considering a significantly high air conditioning load. This condition tends to drive a high load current in all the current carrying equipment causing internal equipment temperature rise. This coupled with the high ambient temperature conditions can lead to failure of multiple components like insulation, adhesives, etc., that can lead to complete equipment failure.
The other aspect that is possible in the field but not reported in the literature is the impact of surrounding equipment heat exhaust causing increased ambient air heating that can cause equipment like switchgear and switchgear panels to be exposed to extreme heat.
Fig. 21 shows the vicinity of the switchgear and switchgear panels around large power transformers. The heat generated within the power transformer is radiated out to cool the transformer, however this causes additional ambient temperature-rise around the transformer where the auxiliary equipment switchgear panels also located.
During some weather conditions heat domes, or severe ice, the switchgear panels have a risk of ice deposition and/or condensation in side the panel enclosures [166]. This may results in a arc flash risk leading to major failures.
Existing Methods to Measure the Impacts: There is some technology available in the market to perform thermal monitoring systems in low and medium voltage switchgear. This temperature monitoring system typically can be used for the detecting irregularities in the following: • The effects of constant thermal cycling on the joints • Corrosion of the metal components damaging insulation • Assets aging, mechanical damage, or environmental factors • Intrusion and human error because of maintenance processes
The key failure modes inside a switchgear are failure of fastened components that can lead to arc flash during extreme heat and icing conditions where the insulation failure is triggered.
Existing Methods to Mitigate the Impacts: The key mitigation methods for outdoor switchgear for extreme heat are Periodic inspections and maintenance; Proper ventilation and if needed forced convection cooling using draft fans, etc.; thermal imaging and alarm systems to detect and manage early failures (22), and; load management which is mostly derating the equipment [167].
@@
For extreme cold operating conditions, understanding and revising the expected extreme cold temperatures and icing conditions followed by conducting the specific type-tests to handle such operating ; Preparation of extreme cold and ice storms includes reinforcements of structures and use of composite insulation with hydrophobic properties. For cabinets, enclosures and panels, use of special gasket materials, anti-condensation heaters to prevent ice build up inside the panels, cabinets and enclosures is needed [169, 170].
Further developments of technologies suggested in the literature include advanced sensors, data driven and physics aware monitoring and mitigation methods [167].
@@ VI-L HVDC Converter Stations
The high voltage direct current (HVDC) Converter stations and associated filter components are a set of infrastructure in themselves [171]. It is divided into Indoor and outdoor components as shown in Fig. 23.
@@
There are other references that show the key components of a HVDC converter station that also illustrate the indoor-outdoor nature of it [173, 174]. There are more than 15 HVDC converter stations in USA with a total capacity greater than 20 GW. The Fig. 23 shows that the main power electronic converters are housed inside a structure called converter Hall that protects the converters from extreme ambient conditions, however, there could be indirect impact of ambient conditions.
Impact of Climate Change: As discussed, parts of a HVDC converter station are outdoor that are directly impacted by the extreme heat, cold and icing conditions. The outdoor components are standard AC equipment and are addressed in the previous section. However, the HVDC converter stations are cooled from external cooling agent like water and in extremely cold conditions and icing conditions, it is important to note that freezing is a common occurrence due to which power electronics can over heat. This failure mode is similar to the failure mode in batteries when the cooling system fails as discussed in subsection V-C. Another failure mode for HVDC systems is also the impact of extreme cold and icing on the HVDC conductors [175] and this is also similar to failure modes for overhead line conductors. There are also indirect impacts of extreme temperatures, i.e., the over loading of the HVDC converter stations and also during heat domes and heat waves the efficiency of cooling systems also goes down that can also lead to indirect failures and possible misoperations of the HVDC converter stations [176]. This can also be a result of extreme temperatures affecting the operations of the reactors and auxiliary equipment present in the HVDC converter stations [176].
Existing Methods to Measure the Impacts: The literature does not report major methods for measuring the impact on extreme temperatures on HVDC converter stations, however since the functioning of the converter stations depends on lot of other substation equipment, it should be taken into account and assessment of all auxiliary equipment in a converter station can help to measure the vulnerability [176].
Existing Methods to Mitigate the Impacts: Apart from other substation equipment vulnerability, the vulnerability of colling systems to extreme cold and icing conditions is important to be addressed. NERC reports on lessons learned [177] and guidelines on winterizing and extreme cold weather preparedness and operations [178] are good references to consider for mitigation of impacts of extreme temperatures. One of the key aspects in the guidelines is recommendations on winterizing liquid cooled systems, ensuring the insulating medium and coolants do not freeze during the extreme cold conditions.
@@ VI-M Power Line Carrier Communication Systems
Power Line carrier communication (PLCC) equipment is typically used for communications over the power carrying conductors. The PLCC system has many different components: coupling capacitor; line trap unit; transmitters and receivers; hybrids and filters; line tuners; master oscillator and amplifiers; protection and earthing of coupling element[179, 180]. The PLC systems have several critical components that need tuning and for accurate functioning, the communication systems have a specific range of operating temperatures. The usual setup of a PLCC system or a power line carrier (PLC) system is illustrated in the Figure below [181].
@@
References [181] indicates several applications of PLCC systems in the future power grid and reference [182] provides the summary of the PLCC market and indicates an increase in the PLCC system market.
Impact of Climate Change: Considering the increased applications listed in [181] and [182] it is important to understand the limitation of the PLCC systems in the future power grid under the consideration of changing climate. While parts of the system are mounted in a control panel either indoors or outdoors, parts of the PLCC system are outdoor and house the coupling circuit and the wave traps. Reference [183] provides details of the PLCC system applications that highlights the indoor and outdoor PLCC systems and provides some details of the ambient operating conditions. These are mainly for the indoor part of the PLCC systems that are usually housed in a panel or enclosure. The key components of a wave trap that is in outdoor equipment that is the Wave Trap or the Line Traps and these are mainly parallel LC circuits (parallel inductor capacitor circuits) in parallel with a lightning arrestor [184, 185]. There is some work done on assessing and considering the impacts of temperatures on wave traps [186]. While reference [186] provides many critical details, it was last updated more than 15 years ago and in light of the recent climate changes, it might have to be reevaluated.
Existing Methods to Measure the Impacts: The key impact of extreme temperature on PLCC equipment is malfunction of the communication equipment. Generally the impact is expected to be temporary like increase in the noise levels, attenuation, however in some rare cases the damage could be permanent and may need replacement of equipment. Reference [187] provides details of monitoring and measurement methods to identify malfunction of various subsystems in a PLCC system.
Existing Methods to Mitigate the Impacts: The key failure mode of PLCC equipment is malfunction of equipment and using the monitoring methods [187] can identify the faulty equipment. The mitigation methods could include having back-up communication methods for essential control applications in the power grid.
@@ VI-N Disconnect Switches and Isolators
Disconnect switches and isolators are electrical switches that control the flow of electricity and isolate circuits when needed for protection.
Impact of Climate Change: Disconnect switches and isolators are susceptible to the changes in operating temperature. This is because the ratings of the switches and isolators are dependent on the ambient temperature. According to IEEE standard C37.30-1997 [188], the usual service conditions are 1) ambient temperatures between -30°C and 40°C, 2) altitude does not exceed 1000 $m$, and 3) wind velocity not exceed 37 $m/sec$. The IEEE standard C37.30-1997 [188] also suggests the unusual service conditions such as 1) higher altitudes will have lower dielectric strength resulting in higher temperature rise in cooling medium (derating/correction factor must be used appropriately), 2) switches within enclosure must use immediate air around the equipment rather than the ambient temperatures, etc.
@@
Existing Methods to Mitigate the Impacts: A method to mitigate the impact on switches and isolators is by installing protection system. However, the protection system’s settings must be adjusted based on the derating behavior of the switches which is not typical.

### v2

@@ VI-I Shunt Reactors
Shunt reactors are important reactive management power devices in the power grid. They are crucial in ensuring that the voltages in the power network do not rise to dangerously high values, causing the insulation in the system to fail. There are both fixed-rating and variable shunt reactors; fixed rating reactors are either dry or oil immersed, and variable reactors are mostly oil immersed. Fixed shunt reactors cannot provide regulation, while variable shunt reactors provide regulation capability (which is necessary with variable renewable generation like wind and solar [155]). Variable shunt reactors have multiple components including a tap changer and cooling system that are vulnerable to extreme temperatures.
Impact of Extreme Temperature Operations: Like oil-cooled transformers, shunt reactors also have temperature-vulnerable components like insulation, windings, tap changer, etc. Reference [156] provides details of the temperature limitations of medium-voltage to high-voltage oil-immersed transformers. Extreme heat (temperatures greater than 40 °C) and extreme cold (temperatures less than $-25$C) ambient conditions are defined as stressed ambient conditions for oil-immersed shunt reactors [156]. There are further details about average temperatures: The monthly average temperature should not exceed 30 °C and annual average should not exceed 20 °C [156]. If the temperatures exceed these ranges, appropriate measures are needed to prevent impacts on the reliability and life of the shunt reactor. These temperatures should be maintained during transport and storage [156]. Reference [157] provides details of mechanical failure and microcracks that can develop in the dry type of shunt reactors. It shows the mechanism of this failure in extreme cold conditions. This can lead to structural failure and eventual fire hazard. This impact is not only the failure of the reactor, but can easily cascade into a fire risk during polar vortex conditions where the wind speeds are relatively high.
Existing Methods to Measure the Impacts: There are not online measurement methods, however Refs. [158] and [159] provide some general guidelines on the performance of the shunt reactors that could be used to assess the impacts of extreme temperatures on the shunt reactors’ performance or the impact of extreme temperatures.
Existing Methods to Mitigate the Impacts: There are not clear insights on the different methods of mitigating the impacts of extreme temperatures on shunt reactors. Because the size and form factor for some of these high-voltage reactors are fairly large, it is a challenging problem to shield these from external extreme ambient temperature conditions.
@@ VI-J SVCs and FACTS Devices
SVC (static VAR compensator) and STATCOM (static synchronous compensator) are shunt devices in the FACTS family. SVCs based on thyristor technology appeared in the field during 1970s. STATCOMs based on GTOs (gate turn-off thyristors) appeared in 1980s, but with the help of high-power IGBTs (insulated gate bipolar transistor) in the 1990s, became commercially available. Both SVC and STATCOM assist in increasing the maximum transfer capability of power lines. This in turn helps with delaying the the need to build new power lines in face of ever-growing power demand on the electric grid [160]. When the space in substations is a constraint. Furthermore, STATCOMs do not have capacitors like SVCs, which comes with a few advantages in terms of vulnerability to extreme temperatures.
SVC and STATCOM offer several benefits: 1) voltage control during normal and contingency conditions, 2) reactive power support during contingencies, 3) voltage stability, 4) overvoltage reduction during lightly loaded conditions, 5) damping oscillations, 6) maximizing power transfer capability, and 7) power quality.
Impact of Extreme Temperature Operations: SVCs have similar points of failures to capacitors and electronics-based devices because of the components within. STATCOMs, which do not have capacitors, have similar but fewer vulnerabilities. Some of the conditions that affect all outdoor equipment are as follows [161], 1) ambient temperature; 2) seismic activity; 3) wind; 4) climatic conditions including high humidity, heavy rainfall or snow, sand-storms, ice load, etc.; and 5) pollution, including dust, corrosive atmosphere, etc.
Existing Methods to Measure the Impacts: Remote monitoring of ambient temperature of the immediate air surrounding the equipment is the main method for assessing potential impacts. [161].
Existing Methods to Mitigate the Impacts: Similarly to frequency changers, these issues can be resolved by placing the equipment in a climate-controlled environment/enclosure. This is especially useful when climate conditions can change drastically at a specific location.
@@ VI-K Switchgear Panels and Control Panels
Switchgear is a class of equipment that houses many different critical components in a substation, switchyard, or other key locations where multiple paths of the power network interface. Indoor switchgear is shielded from the environmental conditions to some extent, but outdoor switchgear is prone to impacts of environmental conditions. In medium-voltage and low-voltage systems, the outdoor switchgear is often metal enclosed. These “weather-proof” housings, however, cannot eliminate the impacts of extreme heat and extreme cold caused by heat domes and polar vortices or winter storms.
There is also outdoor switchgear, typically located in substations, for protecting/managing key substation equipment like transformers and circuit breakers [162]. The radiators and oil pumps are all dedicated low-voltage auxiliary components like motors that are managed with their own switchgear. These are also called the control cabinets, and are mounted either on or near the transformer (21).
@@
Impact of Extreme Temperature Operations: Overheating and insulation failures are the key failure modes for switchgear [163]. The switchgear equipment located in the substation yard is not only subject to weather-related ambient conditions but also the heat generated by neighboring equipment. Especially during extreme heat conditions, the load/demand in the system is high (considering a significantly high air conditioning load). This condition tends to drive a high load current in all the current-carrying equipment, causing internal equipment temperature rise. This, coupled with the high ambient temperature, can lead to failure of multiple components like insulation, adhesives, etc., which in turn can lead to complete equipment failure.
Another aspect that is possible in the field but not reported in the literature is the heat exhaust from surrounding equipment increasing the ambient temperature, which can cause equipment like switchgear and switchgear panels to be exposed to extreme heat.
Fig. 21 shows the vicinity of the switchgear and switchgear panels around large power transformers. The heat generated within the power transformer is radiated out to cool the transformer; however, this causes additional ambient temperature-rise around the transformer where the auxiliary equipment switchgear panels also located.
During some weather conditions the switchgear panels have a risk of ice deposition or condensation inside the panel enclosures [164]. This may result in arc flash risk, leading to major failures.
Existing Methods to Measure the Impacts: There is some technology available in the market to monitor temperatures in low- and medium-voltage switchgear. This temperature monitoring system can be used for detecting irregularities in the following: 1) the effects of constant thermal cycling on the joints; 2) corrosion of the metal components damaging insulation; assets aging, mechanical damage, or environmental factors; and intrusion and human error because of maintenance processes.
The key failure modes of switchgear are failure of fastened components that can lead to arc flash during extreme heat and icing conditions, causing insulation failure.
Existing Methods to Mitigate the Impacts: The key mitigation methods for outdoor switchgear for extreme heat are periodic inspections and maintenance; proper ventilation and (if needed) forced convection cooling using draft fans, etc.; thermal imaging and alarm systems to detect and manage early failures (22); and load management, which is mostly derating the equipment [165].
@@
For extreme cold operating conditions, mitigation measure include understanding and revising the expected extreme cold temperatures and icing conditions followed by conducting the specific type-tests to handle such operation. Preparation for extreme cold and ice storms includes reinforcements of structures and use of composite insulation with hydrophobic properties. For cabinets, enclosures, and panels, the use of special gasket materials, anti-condensation heaters to prevent ice build up inside the panels, cabinets and enclosures is needed [167, 168].
Further developments of technologies suggested in the literature include advanced sensors, data-driven and physics-aware monitoring, and mitigation methods [165].
@@ VI-L HVDC Converter Stations
The high voltage direct current (HVDC) converter stations and associated filter components are a set of infrastructure in themselves [169]. Stations are divided into indoor and outdoor components, as shown in Fig. 23.
@@
Other references that showing the key components of a HVDC converter stations also illustrate their indoor-outdoor nature [171, 172]. There are more than 15 HVDC converter stations in USA with a total capacity greater than 20 GW. Fig. 23 shows that the main power electronic converters are housed inside a structure called converter hall that protects the converters from extreme ambient conditions; however, there could be indirect impact of ambient conditions.
Impact of Extreme Temperature Operations: Outdoor components of an HVDC converter station are directly affected by extreme heat, cold, and icing conditions. The outdoor components are standard AC equipment and are addressed in the previous sections. However, HVDC converter stations are cooled from external cooling agents like water. In extremely cold conditions and icing conditions, it is important to note that freezing is a common occurrence due to which power electronics can over heat. This failure mode is similar to the failure mode in batteries when the cooling system fails, as discussed in subsection V-C. Another failure mode for HVDC systems is also the impact of extreme cold and icing on the HVDC conductors [173] and this is also similar to failure modes for overhead line conductors. There are also indirect impacts of extreme temperatures, i.e., the overloading of the HVDC converter stations. During heat domes and heat waves, the efficiency of cooling systems also goes down, which can also lead to indirect failures and possible misoperations of the HVDC converter stations [174]. This can also be a result of extreme temperatures affecting the operations of the reactors and auxiliary equipment present in the HVDC converter stations [174].
Existing Methods to Measure the Impacts: The literature does not report major methods for measuring the impact of extreme temperatures on HVDC converter stations. However, because the functioning of converter stations depends on the many types of equipment that compose them, assessing all auxiliary equipment in a converter station can help to measure its overall vulnerability [174].
Existing Methods to Mitigate the Impacts: Apart from other substation equipment vulnerabilities, the vulnerability of cooling systems to extreme cold and icing conditions is important to address. NERC reports on lessons learned [175] and guidelines on winterizing and extreme cold weather preparedness and operations [176] are good references to consider for mitigation of impacts of extreme temperatures. One of the key aspects in the guidelines is recommendations on winterizing liquid-cooled systems, ensuring that the insulating media and coolants do not freeze during extreme cold conditions.
@@ VI-M Power Line Carrier Communication Systems
Power line carrier communication (PLCC) equipment is typically used for communications over the power carrying conductors. The PLCC system has many different components: coupling capacitor; line trap unit; transmitters and receivers; hybrids and filters; line tuners; master oscillator and amplifiers; protection and earthing of coupling element[177, 178]. The PLC systems have several critical components that need tuning, and for accurate functioning, the communication systems have a specific range of operating temperatures. The usual setup of a PLCC system or a power line carrier (PLC) system is illustrated in Fig. 24 [179].
@@
Ref. [179] indicates several applications of PLCC systems in the future power grid, and Ref. [180] provides the summary of the PLCC market and indicates an increase in the PLCC system market.
Impact of Extreme Temperature Operations: Considering their increased applications [179, 180], it is important to understand the limitations of the PLCC systems in the future power grid under a changing climate. While parts of the system are mounted in a control panel either indoors or outdoors, parts of the PLCC system are outdoors and house the coupling circuit and the line traps. Reference [181] provides details of the PLCC system applications that highlights the indoor and outdoor PLCC systems and provides some details of the ambient operating conditions. These are mainly for the indoor part of the PLCC systems, which are usually housed in a panel or enclosure. The key components outdoors components of PLCC equipment are the line traps (also called wave traps); these are mainly parallel LC circuits (parallel inductor capacitor circuits) in parallel with a lightning arrestor [182, 183]. There is some work done on assessing and considering the impacts of temperatures on line traps [184]. While reference [184] provides many critical details, it was last updated more than 15 years ago. It may need to be reevaluated in light of recent climate changes.
Existing Methods to Measure the Impacts: The key impact of extreme temperature on PLCC equipment is malfunction of the communication equipment. Generally, the impact is expected to be temporary and includes increase in the noise levels and attenuation. However, in some rare cases, the damage could be permanent and require replacement of equipment. Reference [185] provides details of monitoring and measurement methods to identify malfunctions of various subsystems in a PLCC system.
Existing Methods to Mitigate the Impacts: The key failure mode of PLCC equipment is malfunction of equipment, and using the monitoring methods [185] can identify the faulty equipment. The mitigation methods could include having back-up communication methods for essential control applications in the power grid.
@@ VI-N Disconnect Switches and Isolators
Disconnect switches and isolators are electrical switches that control the flow of electricity and isolate circuits when needed for maintenance.
Impact of Extreme Temperature Operations : Disconnect switches and isolators are susceptible to the changes in operating temperature. This is because the ratings of the switches and isolators are dependent on the ambient temperature. According to IEEE standard C37.30-1997 [186], the usual service conditions are 1) ambient temperatures between -30°C and 40°C, 2) altitude does not exceed 1000 $m$, and 3) wind velocity not exceed 37 $m/sec$. The IEEE standard C37.30-1997 [186] also suggests the unusual service conditions such as 1) higher altitudes will have lower dielectric strength resulting in higher temperature rise in cooling medium (derating/correction factor must be used appropriately), 2) switches within enclosure must use immediate air around the equipment rather than the ambient temperatures, etc.
@@
In case of extreme cold, the switches has to pass the ice breaking test which states that the rated ice-breaking ability are 9.5 mm and 19 mm [188]. The rated ice-breaking ability is the maximum thickness of ice deposited on a device that will not interfere with its successful opening or closing. Under certain circumstances, an ice storm can cause a deposit of ice of such thickness that overhead lines fail and operation of specific switching equipment is impaired [127]. Ice is produced naturally in two ways, they are 1) clear ice appears from rain falling through air with a temperature between 0°C (+32F) and –10°C (+14F)., and 2) Rime ice, characterized by a white appearance from the air entrapped during ice formation, appears from rain falling through air with a temperature below –10°C (+14 °F) or from condensation of atmospheric moisture on cold surfaces.
@@
Existing Methods to Mitigate the Impacts: A method to mitigate the impact on switches and isolators is by installing protection system (for motorized switches). However, the protection system’s settings must be adjusted based on the derating behavior of the switches which is not typical.
VI-O Backup Batteries
Most of the substation equipment like relays, motors for circuit breakers, re-closers, isolators, sectionalizers, etc., need electric power supply to operate. Most substations have a dedicated backup power supply in the form of batteries that are mainly used to support the substation operation including the relay and other critical equipment that is housed inside the substation control room.
Impact of Extreme Temperature Operations: The battery systems in the substations are usually stored in a separate battery room and this room needs to be maintained at a temperature of 25°$\degree$°C. Else the battery life is significantly shortened. Extreme heat causes a shortening of the lifespan of the battery. Extreme cold causes the a reduced charge and increased Sulfation leading to corrosion and overall deterioration of the battery [189]. Some specifications show that life of a Ni-Cad battery is better compared to a lead-acid battery with increased temperatures [189]. Extreme heat can cause irreversible damage to the battery systems [190].
Existing Methods to Measure the Impacts: There are some devices available to measure the health of a stationary battery system like the ones used for backup power in substations called battery testers [191]. While these battery testers are well developed for lead-acid battery chemistry. There are other methods that are more suitable for other battery chemistry that are described in detail in reference [192], where there are methods and metrics that are evaluated to measure the health of a battery system. Since extreme cold and extreme heat operations of the battery impact the state of the charge, battery end-of-life (EOL), the methods described in reference [192] can be directly applied.
Existing Methods to Mitigate the Impacts: The key methods for mitigating the adverse impacts of extreme temperature operation are to have a temperature controlled environment of the batteries. This is however a challenge in peak summer and peak winter durations, especially under events like heat domes and polar vortices as the external temperature often tends to limit the amount of cooling/heating a heating ventilating and air conditioning (HVAC) system could provide [193]. ANSI/IEEE 484 provides more details of the battery room requirements [194]. Reference [195] provide details of operating temperatures and other considerations for designing and operating battery rooms in a substation.